# MnLargeSymbols’164 MnLargeSymbols’171 MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ 3.2 Adaptive Batch Size Schedules with Data Parallelism
It is worth noting that efficiently implementing the approximate norm test (3) in deep learning libraries such as PyTorch (Paszke et al., 2019) is highly nontrivial, since per-sample gradients $\nabla\ell_{i}(w_{k})$ are unavailable in the backward step of a standard training loop, but only the batch gradient $\nabla\mathscr{L}_{\mathcal{B}_{k}}(w_{k})$ under a single-device setting or the minibatch gradient $\nabla\mathscr{L}_{\mathcal{B}_{k,j}}(w_{k})$ at each worker $j$ under PyTorch DDP. If we were to implement the native approximate norm test (3), we would have had to compute per-sample gradients in parallel using vectorized mappings and based on a deep copy of the model, leading to undesirable memory and computational overheads. Thus, in practical implementation under data parallelism, instead of the approximate norm test (3), we propose to make use of the minibatch gradients of the workers to construct an estimator for the gradient variance $\widehat{\mathrm{Var}}_{i\in\mathcal{B}_{k}}(\nabla\ell_{i}(w_{k}))\coloneqq\frac{1}{J}\sum_{j\in\llbracket J\rrbracket}\left(\nabla\mathscr{L}_{\mathcal{B}_{k,j}}(w_{k})-\nabla\mathscr{L}_{\mathcal{B}_{k}}(w_{k})\right)^{2}$, leading to the following more efficient implementation:
@@ 3.3 Adaptive Batch Size Schedules with 2D Parallelism via PyTorch FSDP
To enable the training of models with more than billions of parameters, model-parallel training presents a more sophisticated paradigm of parallelism. It shards the parameters of models and allocates different shards to different workers. In essence, PyTorch FSDP (Zhao et al., 2023), which shares similarities with ZeRO-3 (Rajbhandari et al., 2020; Ren et al., 2021) in DeepSpeed (Rasley et al., 2020), operates by substituting the all-reduce operation in PyTorch DDP with all-gather and reduce-scatter operations.
@@ 4 Convergence Analysis
Complementary to the convergence results of the norm test for SGD (De et al., 2016, 2017) and AdaGrad (Lau et al., 2024b), we derive convergence guarantees for Adam, acknowledging its prevalence in training deep neural networks for both computer vision and, more recently, language models. Adam (Kingma and Ba, 2015) employs the following update formula:
@@
Similarly to the analysis for AdaGrad (Lau et al., 2024b), we also require a coordinate-wise version of the (exact variance) norm test to hold due to the use of adaptive step size.
@@
Following closely a similar analysis to that in Wang et al. (2023), we provide the following convergence results of the norm test for Adam.
@@
The full statement of this theorem and its proofs, as well as more in-depth related discussions, are deferred to Appendix B. The convergence results presented do not account for the decoupled weight decay in AdamW (Loshchilov and Hutter, 2019), which is more commonly used as an optimizer for language model pretraining. Furthermore, considerations such as learning rate schedules and gradient clipping are not included in these findings. Extending the above convergence guarantees to these settings is highly challenging and nontrivial and is left for future work.
@@ 5 Numerical Experiments
To showcase the versatility and scalability of FSDP-Norm, we conduct experiments with various families of decoder-only autoregressive language models at with different sizes and pretraining datasets. These include MicroLlama 300M (Wang, 2024), TinyLlama 1.1B (Zhang et al., 2024b) and OpenLlama 3B (Geng and Liu, 2023) on the C4 dataset (Raffel et al., 2020). The C4 dataset are tokenized using the Llama 2 tokenizer (Touvron et al., 2023) with a vocabulary size of 32,000. Experiments are conducted on workstations equipped with 4 NVIDIA L40S GPUs (MicroLlama) and 4 NVIDIA A100-SXM 80GB GPUs (TinyLlama and OpenLlama). The training of the latter two models only feasible with PyTorch FSDP but not with PyTorch DDP using such hardware configurations, even with mixed-precision training (bfloat16 is used). Our implementation utilizes the PyTorch FSDP API in PyTorch 2.6.1 and is simplified through Lightning Fabric of Lightning 2.4 (Falcon and The PyTorch Lightning team, 2019). For the ease of training language models, we also use LitGPT 0.5.3 (Lightning AI, 2023). Open-source implementation of DDP-Norm and FSDP-Norm will be made available as a GitHub repository upon the acceptance of this manuscript.
@@
Adhering to the pretraining configurations of open-source LLMs such as TinyLlama (Zhang et al., 2024b) and OLMo (Groeneveld et al., 2024), our training specifications include a linear warmup followed by a cosine decay learning rate schedule, and the AdamW optimizer with weight decay and gradient clipping. The adaptive batch size schedule is set to a maximum global batch size, above which the norm test is no longer performed, opting for fixed interval testing over step-by-step (a test interval 1 is used, but longer interval entails reduced overheads brought by the test). Efficiency dictates using the test in its original form rather than its coordinate-wise variant, despite convergence guarantees. Given that batch sizes increase to the maximum possible values in the early stages, we only pretrain our models for a number of samples that are sufficient to display the behavior of our method, treating these experiments mainly as proofs of concept. Detailed configurations are provided in Appendix C.
@@ 5.1 MicroLlama 300M
We first pretrain MicroLlama with 300M trainable parameters on the C4 dataset (Raffel et al., 2020) under the same sets of other hyperparameters in order to better understand the effect of adaptive batch sizes. We compare with various constant batch size baselines $b_{k}\in\{2048,4096,8192\}$ and a stagewise batch size schedule 2048-4096-8192 for 2.5-2.5-95% of training tokens mimicking a popular batch size warmup for pretraining LLMs, and plot the results in Figure 2. We apply DDP-Norm for this relatively small model to demonstrate the applicability of the proposed schedules with PyTorch DDP. In Table 1, we report the total number of gradient steps (step), average batch size (bsz.), wall-clock time (time; in hours), best training loss (loss) and best validation loss (val loss; estimated by 100 iterations).
@@ 5.4 Further Discussions of Experimental Results
The hyperparameter $\eta$ in the adaptive batch size schedules has the effect of controlling the probability of obtaining a descent direction and hence increasing the batch size. Obviously, choosing a right value of $\eta$ is vital for our method to succeed. Across all three sets of experiments of different model scales, we found that larger values of $\eta$ generally lead to more gradual batch size increments, but smaller values would allow full utilization of available compute resources at earlier stages of training but might defeat the prupose of adaptive batch sizes. Note that $\eta$ also varies with the base learning rate $\alpha$ and the quality of the training datasets. In the series of works of adaptive sampling methods (Byrd et al., 2012; Bollapragada et al., 2018; Bollapragada and Wild, 2023), there are in-depth discussions on choosing the learning rate via some line-search procedures, which are however usually infeasible when training large deep neural networks.
@@
We conjecture that there are more general scaling laws of the critical batch size (see e.g., Kaplan et al., 2020; Su et al., 2024; Sclocchi and Wyart, 2024; Zhang et al., 2024a) in relation to $\eta$ which controls gradient approximation quality and the scale of gradient noise. For most choices of $\eta$ in the three sets of experiments, we choose $\eta$ small enough so that global batch sizes increase rapidly and reach the maximum possible values. However, in Figure 2, when $\eta=0.15$, the final batch size is around 3800, which might be the critical batch size at this value of $\eta$. It is thus crucial to understand the notion of critical batch sizes through the lens of gradient approximation quality and we leave this for future work.
@@ 6 Concluding Remarks
We create an efficient PyTorch FSDP implementation of the norm test for large-scale distributed training, focusing on hardware use and ease of development. Our implementation shows that adaptive batch size schedules can pretrain Llama 2 language models with up to 3 billion parameters using few GPUs (Touvron et al., 2023). Furthermore, we provide convergence guarantees of the norm test for Adam, suggesting that our proposed adaptive batch size schedules are not only practically feasible, but also theoretically principled. Due to its generality, versatility, and scalability, we foresee extensive use of the adaptive batch size schedules in pretraining large transformer models like vision transformers (ViT; Dosovitskiy et al., 2021) and autoregressive image models (Aim; El-Nouby et al., 2024). We emphasize our attention on a PyTorch FSDP approach due to its integration with PyTorch. However, a more advanced implementation of the adaptive batch size schedules, using a new version of PyTorch FSDP (FSDP2) and tensor parallelism via PyTorch DTensor (Distributed Tensor), will significantly enhance the scalability of the method for training models exceeding 7B parameters with 2D or 3D parallelism. For further exploration, refer to the torchtitan (Liang et al., 2024) and lingua (Videau et al., 2024) repositories.
@@ Appendix B Proofs of Main Text
We give a brief sketch of the omitted proof of the main text in this section. Notice that the convergence analysis of the norm test for Adam largely follows that in Wang et al. (2023), where more details and remarks of the analysis and rationales of its derivation can be found.
@@
Despite the similarity of the proof techniques, we emphasize that our setting requires less restrictive assumptions. While Wang et al. (2023) assume the stochastic oracle of the gradient (i.e., batch gradient in our case) has coordinate-wise affine variance, i.e., for any batch of samples $\mathcal{B}\subseteq\mathcal{D}_{n}$ and $(\sigma,\tau)\in\left(0,\infty\right)^{2}$, we have
@@
which is also known as the coordinate-wise expected strong growth (E-SG) condition (Lau et al., 2024b). Note that the coordinate-wise (E-SG) condition implies the coordinate-wise relaxed growth (RG) condition (Bottou et al., 2018), adopting the nomenclature in Khaled and Richtárik (2023):
@@ B.1 Technical Lemmas
We state without proof the following technical lemmas from Wang et al. (2023).
@@ B.2 Proof of Theorem 1
Since the proof of Theorem 1 is highly similar to that in Wang et al. (2023), we just provide a proof sketch. We state the formal theorem as follows.
@@ Appendix C Details of Numerical Experiments
We provide a summary table for the architecture of the language models we pretrained. More details of these models can be found at Wang (2024); Zhang et al. (2024b); Geng and Liu (2023).

### v2

@@ 3.2 Adaptive Batch Size Schedules with Data Parallelism
It is worth noting that efficiently implementing the approximate norm test (3) in deep learning libraries such as PyTorch [57] is highly nontrivial, since per-sample gradients $\nabla\ell_{i}(w_{k})$ are unavailable in the backward step of a standard training loop, but only the batch gradient $\nabla\mathscr{L}_{\mathcal{B}_{k}}(w_{k})$ under a single-device setting or the minibatch gradient $\nabla\mathscr{L}_{\mathcal{B}_{k,j}}(w_{k})$ at each worker $j$ under PyTorch DDP. If we were to implement the native approximate norm test (3), we would have had to compute per-sample gradients in parallel using vectorized mappings and based on a deep copy of the model, leading to undesirable memory and computational overheads. Thus, in practical implementation under data parallelism, instead of the approximate norm test (3), we propose to make use of the minibatch gradients of the workers to construct an estimator for the gradient variance
$\widehat{\mathrm{Var}}_{i\in\mathcal{B}_{k}}(\nabla\ell_{i}(w_{k}))\coloneqq\frac{1}{J}\sum_{j\in\llbracket J\rrbracket}\left(\nabla\mathscr{L}_{\mathcal{B}_{k,j}}(w_{k})-\nabla\mathscr{L}_{\mathcal{B}_{k}}(w_{k})\right)^{2},$
leading to the following more efficient implementation:
@@ 3.3 Adaptive Batch Size Schedules with 2D Parallelism via PyTorch FSDP
To enable the training of models with more than billions of parameters, model-parallel training presents a more sophisticated paradigm of parallelism. It shards the parameters of models and allocates different shards to different workers. In essence, PyTorch FSDP [88], which shares similarities with ZeRO-3 [61, 63] in DeepSpeed [62], operates by substituting the all-reduce operation in PyTorch DDP with all-gather and reduce-scatter operations.
@@ 4 Convergence Analysis
Complementary to the convergence results of the norm test for SGD [13, 14] and AdaGrad [42], we derive convergence guarantees for Adam, acknowledging its prevalence in training deep neural networks for both computer vision and, more recently, language models. Adam [36] employs the following update formula (with bias corrections for $m_{k}$ and $v_{k}$ dropped):
@@
Similarly to the analysis for AdaGrad [42], we also require a coordinate-wise version of the (exact variance) norm test to hold due to the use of adaptive step size.
@@
Following closely a similar analysis to that in [76], we provide the following convergence results of the norm test for Adam.
@@
The full statement of this theorem and its proofs, as well as more in-depth related discussions, are deferred to Appendix B. The convergence results presented do not account for the decoupled weight decay in AdamW [51], which is more commonly used as an optimizer for language model pretraining. Furthermore, considerations such as learning rate schedules and gradient clipping are not included in these findings. Extending the above convergence guarantees to these settings is highly challenging and nontrivial and is left for future work.
@@ 5 Numerical Experiments
To showcase the versatility and scalability of FSDP-Norm, we conduct experiments with various families of decoder-only autoregressive language models at with different sizes and pretraining datasets. These include MicroLlama 300M [77], TinyLlama 1.1B [84] and OpenLlama 3B [25] on the C4 dataset [60]. The C4 dataset are tokenized using the Llama 2 tokenizer [73] with a vocabulary size of 32,000. Experiments are conducted on workstations equipped with 4 NVIDIA L40S GPUs (MicroLlama) and 4 NVIDIA A100-SXM 80GB GPUs (TinyLlama and OpenLlama). The training of the latter two models only feasible with PyTorch FSDP but not with PyTorch DDP using such hardware configurations, even with mixed-precision training (bfloat16 is used). Our implementation utilizes the PyTorch FSDP API in PyTorch 2.6.1 and is simplified through Lightning Fabric of Lightning 2.4 [20]. For the ease of training language models, we also use LitGPT 0.5.3 [48]. Open-source implementation of DDP-Norm and FSDP-Norm is available at https://github.com/timlautk/adaptive-batch-fsdp.
@@
Adhering to the pretraining configurations of open-source LLMs such as TinyLlama [84] and OLMo [27], our training specifications include a linear warmup followed by a cosine decay learning rate schedule, and the AdamW optimizer with weight decay and gradient clipping. The adaptive batch size schedule is set to a maximum global batch size, above which the norm test is no longer performed, opting for fixed interval testing over step-by-step (a test interval 1 is used, but longer interval entails reduced overheads brought by the test). Efficiency dictates using the test in its original form rather than its coordinate-wise variant, despite convergence guarantees. Given that batch sizes increase to the maximum possible values in the early stages, we only pretrain our models for a number of samples that are sufficient to display the behavior of our method, treating these experiments mainly as proofs of concept. Detailed configurations are provided in Appendix C.
@@ 5.1 MicroLlama 300M
We first pretrain MicroLlama with 300M trainable parameters on the C4 dataset [60] under the same sets of other hyperparameters in order to better understand the effect of adaptive batch sizes. We compare with various constant batch size baselines $b_{k}\in\{2048,4096,8192\}$ and a stagewise batch size schedule 2048-4096-8192 for 2.5-2.5-95% of training tokens mimicking a popular batch size warmup for pretraining LLMs, and plot the results in Figure 2. We apply DDP-Norm for this relatively small model to demonstrate the applicability of the proposed schedules with PyTorch DDP. In Table 1, we report the total number of gradient steps (step), average batch size (bsz.), wall-clock time (time; in hours), best training loss (loss) and best validation loss (val loss; estimated by 100 iterations).
@@ 5.4 Further Discussions of Experimental Results
The hyperparameter $\eta$ in the adaptive batch size schedules has the effect of controlling the probability of obtaining a descent direction and hence increasing the batch size. Obviously, choosing a right value of $\eta$ is vital for our method to succeed. Across all three sets of experiments of different model scales, we found that larger values of $\eta$ generally lead to more gradual batch size increments, but smaller values would allow full utilization of available compute resources at earlier stages of training but might defeat the prupose of adaptive batch sizes. Note that $\eta$ also varies with the base learning rate $\alpha$ and the quality of the training datasets. In the series of works of adaptive sampling methods [12, 8, 7], there are in-depth discussions on choosing the learning rate via some line-search procedures, which are however usually infeasible when training large deep neural networks.
@@
We conjecture that there are more general scaling laws of the critical batch size (see e.g., [33, 70, 64, 83]) in relation to $\eta$ which controls gradient approximation quality and the scale of gradient noise. For most choices of $\eta$ in the three sets of experiments, we choose $\eta$ small enough so that global batch sizes increase rapidly and reach the maximum possible values. However, in Figure 2, when $\eta=0.15$, the final batch size is around 3800, which might be the critical batch size at this value of $\eta$. It is thus crucial to understand the notion of critical batch sizes through the lens of gradient approximation quality and we leave this for future work.
@@ 6 Concluding Remarks
We create an efficient PyTorch FSDP implementation of the norm test for large-scale distributed training, focusing on hardware use and ease of development. Our implementation shows that adaptive batch size schedules can pretrain Llama 2 language models with up to 3 billion parameters using few GPUs [73]. Furthermore, we provide convergence guarantees of the norm test for Adam, suggesting that our proposed adaptive batch size schedules are not only practically feasible, but also theoretically principled. Due to its generality, versatility, and scalability, we foresee extensive use of the adaptive batch size schedules in pretraining large transformer models like vision transformers (ViT) [18] and autoregressive image models (Aim) [19]. We emphasize our attention on a PyTorch FSDP approach due to its integration with PyTorch. However, a more advanced implementation of the adaptive batch size schedules, using a new version of PyTorch FSDP (FSDP2) and tensor parallelism via PyTorch DTensor (Distributed Tensor), as well as availing of stronger computational hardware, will significantly enhance the scalability of the method for training models exceeding 7B parameters with 2D, 3D or even 4D parallelism. For further exploration, we refer readers to the torchtitan [47] and lingua [75] repositories. Furthermore, while our current implementation is based on PyTorch FSDP, but is readily extendable to other deep learning frameworks such as JAX [10] with FSDP and/or GShard [44].
Limitations.
In this work, we are primarily concerned with model generalization performance measured by validation loss without any evaluation on downstream benchmarks. The main reason for this is that we did not fully pretrain the models for sufficient number of tokens, implying that these models will not be competitive on downstream benchmarks. However, we expect that models fully pretrained with our proposed schedules will achieve very competitive performance on the evaluation of downstream benchmarks. We also remark that we can also incorporate other paradigms of parallelism such as pipeline and context parallelism with our proposed scheme, leading to 4D parallelism (data, tensor, pipeline, context parallel) for large-scale pretraining. While not supported in our current implementation, this can be achieved using the recent library picotron [87] or the more sophisticated library Megatron-LM [66]. We leave this implementation for future work.
@@ Appendix B Proofs of Main Text
We give a brief sketch of the omitted proof of the main text in this section. Notice that the convergence analysis of the norm test for Adam largely follows that in [76], where more details and remarks of the analysis and rationales of its derivation can be found.
@@
Despite the similarity of the proof techniques, we emphasize that our setting requires less restrictive assumptions. While Wang et al. [76] assume the stochastic oracle of the gradient (i.e., batch gradient in our case) has coordinate-wise affine variance, i.e., for any batch of samples $\mathcal{B}\subseteq\mathcal{D}_{n}$ and $(\sigma,\tau)\in\left(0,\infty\right)^{2}$, we have
@@
which is also known as the coordinate-wise expected strong growth (E-SG) condition [42]. Note that the coordinate-wise (E-SG) condition implies the coordinate-wise relaxed growth (RG) condition [9], adopting the nomenclature in [35]:
@@ B.1 Technical Lemmas
We state without proof the following technical lemmas from [76].
@@ B.2 Proof of Theorem 1
Since the proof of Theorem 1 is highly similar to that in [76], we just provide a proof sketch. We state the formal theorem as follows.
@@ Appendix C Details of Numerical Experiments
We provide a summary table for the architecture of the language models we pretrained. More details of these models can be found at [77, 84, 25].